OpenStudy (anonymous):

Solve, assuming a linear equation fits the situation. The value of a graphing calculator is $225. After 2 years, the value of this calculator is $160. Find the value of the calculator after 5 years. ( Hint, Let the time now be 0, the time in two years be 2, and the time in five years be 5.)

OpenStudy (anonymous):

Let x be the number of years from now. Value of calculator = 225- x(c), where c is a constant. When x=2, Value of calculator = 160 = 225 - 2C 2c=65 C= 32.5 When x=5, Value of calculator = 225 - 5(32.5) =62.5

OpenStudy (campbell_st):

you have the data points (0, 225) and (2, 160) find the gradient of the line m = (225 - 160)/(0 - 2) 65/-2 = -32.5 next use the point slope formula to find the equation of the line \[y - y _{1} = m( x -x _{1})\] you have m = 32.5 and x1 = 0 and y1 = 225 substitute and simplfy to find the equation of the line

OpenStudy (anonymous):

So whats the final answer in less confusing terms

OpenStudy (campbell_st):

well you have y - 225 = -32.5(x - 0) that should make it easy for you to simplify

OpenStudy (campbell_st):

oops the gradient is -32.5

OpenStudy (anonymous):

Can you just tell me the answer im tired and have no calculator @campbell_st

OpenStudy (campbell_st):

you don't need a calculator... and its very easy you need the equation in the form y = mx + b

OpenStudy (anonymous):

Can you just tell me what it is for just this one time?

OpenStudy (campbell_st):

sorry... its rather straight forward

OpenStudy (anonymous):

Well im sorry but i dont understand it??

OpenStudy (anonymous):

And if its that straight forward as you put it why cant u juss tell me

OpenStudy (campbell_st):

well perhaps more revision is needed

OpenStudy (anonymous):

I just dont understand cant you just tell me what the answer is just this once!!

OpenStudy (anonymous):

@campbell_st WOW you cant seriously just tell me the answer one time??! im sure anyone else would for you if you needed help

OpenStudy (campbell_st):

here's a clue...b = 225

OpenStudy (anonymous):

Thats the dumbest clue ever!! i basically just told you that in the beginning butt wipe

OpenStudy (campbell_st):

well you know b and I've calculated m... and I've written the point slope form of the line... just substitute

OpenStudy (anonymous):

Didnt want to intrude in this argument. But i tot it was getting out of hand. I gave the answer already. BUt campbell gave a better explanation. And i side campbell, because the calculation is indeed not very difficult. Even if you do not have a calculator, taking the time to type the numbers in Google, will pretty much give you the answer.